OpenStudy (anonymous):

Perform the Indicated operation (x+1)^3 in Z3[x]

OpenStudy (anonymous):

\[ (x+1)^{3} in Z_{3}[x]\]

terenzreignz (terenzreignz):

\[\huge (x+1)^3 \ \ \ \ in \ \mathbb{Z}_3[x]\] LOL any idea how to do this? :)

OpenStudy (anonymous):

do i just use f(0) and f(1) to solve for them in Z3? I'm not sure really thats why I'm asking and my book doesn't even address it.

OpenStudy (anonymous):

hint \(3\equiv 0\)

terenzreignz (terenzreignz):

Freshman's dream :D Anyway, evaluate this (x+1)^3 normally first, and tell me what you get.

OpenStudy (anonymous):

lol

terenzreignz (terenzreignz):

Cube of a binomial?

OpenStudy (anonymous):

1+3 x+3 x^2+x^3

terenzreignz (terenzreignz):

That's right :) \[\huge x^3 + 3x^2+3x+1\]but... we take everything in mod 3... In particular, in \[\large \mathbb{Z}_3\] 3=0, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so is it X^3 + 1?

terenzreignz (terenzreignz):

You got it :) And now, for any prime number p

terenzreignz (terenzreignz):

\[\huge (a+b)^p=a^p+b^p \ \ \ \ \ in \ \mathbb{Z}_p\]

OpenStudy (anonymous):

i always suspected my math teachers were lying, and \((a+b)^3=a^3+b^3\)

terenzreignz (terenzreignz):

Oh abstract algebra... so deceptive :D That was fun... but I'm getting drowsy :/ ----------------------------------------- Terence out

OpenStudy (anonymous):

Thank you